# In-vitro measurements coupled with in-silico simulations for stochastic calibration and uncertainty quantification of the mechanical response of biological materials

### v1

@@ 3.1 Bayesian optimization
We first performed specimen-specific parameter calibration using Bayesian optimization (BO), and the results observed for a typical specimen are represented in Figures 4, 5, 6, and 7. Figure 4 shows that the minimum $f(x)$ function converges rapidly, even with only a few initial iterations. This characteristic makes BO popular in machine learning for hyperparameter optimization, where the learning algorithm is treated as a black-box function. Consider the convergence plot given in Figure 4, although the initial minimum value $f(x)$ is 65 N, after 100 iterations (or function calls), it decreases to less than 20 N. In addition, Figure 5 presents the in-vitro measurements obtained from experiments and the in-silico observations generated using optimized FRVE parameter sets in BO. Here, we note that the terms ‘ground truth’, ‘in-vitro measurements’, and ‘experimental observations’ are often used interchangeably to convey a similar meaning. Furthermore, ‘in-silico simulation’, ‘black-box simulator’, and ‘numerical observations’ are also often used interchangeably to refer to a similar meaning. Figure 5 shows that the minimum $f(x)$ values of the objective function, determined by incremental cyclic loading test measurements in one step, two steps, and incremental steps, are 5.693 N, 3.871 N, and 13.232 N, respectively. These results show that the FRVE model can capture the mechanical response of the tendon. Furthermore, we have also included the updated parameter set that was sampled during BO for each iteration as additional figures in this study (See supplementary Graphs G1).
We performed the BO process for 100 calls of the black-box simulator, and the parameter sensitivity can probably be improved once the iteration number is increased. However, this can increase computational time, and we have not focused on characterizing the parameters with a high rank of sensitivity. The target of using BO is to determine how the specimen-specific parameters vary across specimens and to use these parameters to predict the priors for ABC inference. We present some parameters for matrix and fiber constituent pairs that were sampled during BO updating as scatter (Figure 6) and contour (Figure 7) plots to visualize where they are mainly clustered. For example, consider the sample points of $E_{m}$ and $E_{f}$; the elastic modulus of the matrix and fiber components is predominantly updated around 60 MPa and 240 MPa, respectively. The remaining parameters for a typical specimen-specific calibration are given as additional figures (See Supplementary Figure S1 and S2). We then automated the process to characterize the remaining 26 samples, and the overall results are given in Table 2. These results suggest that inter-subject or inter-specimen variability and uncertainty in the biomechanical properties of these materials are high.
@@
Site ID $\tau_{\text{1m}}$ (s) $\tau_{{2\text{m}}}$ (s) $\tau_{{3\text{m}}}$ (s) g1m (MPa) g2m (MPa) g3m (MPa) Em (MPa) vm $\tau_{\text{1f}}$ (s) $\tau_{\text{2f}}$ (s) $\tau_{\text{3f}}$ (s) g1f (MPa) g2f (MPa) g3f (MPa) Ef (MPa) $\alpha$f RMSE (N)
@@
For the clinical aspect, correlation analysis between micro-scale matrix and fiber parameters calibrated via BO can also be performed. Here, to account for dependency due to repeated observations for the same patient’s specimen harvested from distal, middle, and proximal sites, we used the rm_corr() repeated measures correlation function from the Pingouin statistical package. Repeated measures reveal that the most significant correlations are observed between $\tau_{3m}$ and $\tau_{1f}$ (r = 0.8), $g_{3m}$ and $g_{3f}$ (r = 0.71), $\tau_{2m}$ and $\tau_{1f}$ (r = 0.69), $g_{2m}$ and $\tau_{1f}$ (r = 0.65), $g_{2m}$ and $g_{3f}$ (r = 0.64), while the only negative significant correlation exists between $E_{m}$ and $\tau_{1f}$ (r = 0.48) (p $<$ 0.05) (Figure 8).
Figure 8: Correlation levels between matrix and fiber variables. *Repeated measures correlation coefficient (r) is significant at the p $<$ 0.05 level
@@ 3.2 Sensitivity analysis
To determine the influence of each parameter on the objective function of FRVE model parameters, we investigate the percentage importance of the parameters using random forest and their sensitivity using Sobol. Here, we use the error function observations $\min f(x)$ given in Eq. 9 as outputs during sensitivity analysis and importance identification. The results of the random forest for a typical specimen in three loading conditions are represented in Figure 9. As expected, the elastic modulus of the fiber $E_{f}$, the exponential coefficient $\alpha_{f}$, and the elastic modulus of the matrix $E_{m}$ have the highest feature importance. Interestingly, the viscoelastic characteristics $g_{im}$ and relaxation time coefficients $\tau_{im}$ of the ground matrix at the $i$th Prony order are more important than the fiber’s parameters for the combined three-loading $\min f(x)$ function.
@@
For Sobol’s sensitivity analysis, the parameter space is created using Saltelli sampling. We generate 20,000 input sampling sets, each consisting of 16 parameters (8 for the ground matrix and 8 for the fiber), and collect the output $\min f(x)$ observations. Similar to the results from the random forest, $E_{f}$, $\alpha_{f}$, $g_{1m}$, and $E_{m}$ are the most sensitive parameters in the model. The total and first-order indices observed for the eight most sensitive parameters are given in Figure 10. Considering total-order index scores, the sequence from highest to lowest is as follows: $E_{f}$ (0.51), $\alpha_{f}$ (0.48), $E_{m}$ (0.45), $g_{1m}$ (0.36), $\tau_{1m}$ (0.34), $v_{m}$ (0.25), and $\tau_{1f}$ (0.22). These results help to focus on the most critical parameters when defining priors for simulation-based ABC inference.
@@ 3.3 Approximate Bayesian computation
We presented specimen-specific parameter identification results in the previous section and observed that the variability between samples is high and may be influenced by many factors, such as the donor’s age, genetics, the tissue’s pathological state, gender, and other specific factors related to tissue characteristics. To determine the uncertainty in these parameters and stochastically calibrate them, we implement likelihood-free inference in a probabilistic framework. The ABC inference is performed with an initial tolerance threshold of $\epsilon=0.3$ and a posterior sample of 40. The inferred parameters for the distal, middle, and proximal sites are given in Table 3. Consider $E_{f}$ at the distal site: the mean is 216.2 MPa, and the 95% HDI bounds are 164.6 MPa and 267.8 MPa.
@@
We have visualized the posterior joint distributions to investigate the relationships, correlations, and uncertainty between the parameters after updating beliefs based on in vitro measurement observations. Some parameter pairs plotted for the matrix and fiber components, as well as their distributions for the distal site, are shown in Figure 11. Consider $g_{1}$: there is a significant negative correlation between the matrix and the fiber component (r = -0.37, p $<$ 0.05). However, the remaining parameters shown in Figure 11 for the matrix ($\tau_{1m}$, $g_{2m}$, $E_{m}$, $v_{m}$) and the fiber ($\tau_{1f}$, $g_{1f}$, $g_{2f}$, $\alpha_{f}$) are not significantly correlated. The joint posterior plots inferred for 16 parameters for three sites are given in Supplementary Figures S3-S5. Once the calibration of the parameters is complete and their uncertain ranges for the highest density interval (HDI) 95% are inferred, the next step is to investigate how well the simulated data from the posterior $f(y_{sim}|\theta)$ can capture the uncertainties of in vitro observations $y_{obs}$. We perform forward simulations based on posterior parameters ($\theta$) for two cases. In the first case, we performed three simulations: one based on the mean and the other two based on the low and high bounds of $\theta$ given in Table 3. We then collected observations from $y_{sim}$. The inferred observation band is shown in cyan, and its bounds are represented in blue in Figure 12. In the second case, we performed 40 simulations, each corresponding to a single posterior parameter set, and then computed the 95% confidence interval (CI) of the 40 $y_{sim}$ observations. The observation band for the 95% CI and its bounds are represented in red (Figure 12). The results obtained in both cases demonstrate that the stochastically calibrated model parameters can capture the uncertainty band of in vitro measurements ($y_{obs}$). However, to characterize and simulate the soft tissue response under extreme uncertainty, it is preferable to use the 95% HDI bound of $\theta$, as demonstrated in case 1.
@@
In some clinical biomechanical studies, it may be necessary to compare the data obtained from different groups. This is usually carried out by applying parametric or non-parametric statistical tests to the results. This approach is effective within a frequentist framework. Although this study does not directly investigate the differences between the three sites, we performed unsupervised machine learning to determine whether the probabilistic data of the groups are clustered and separated. Here, we primarily used principal component analysis (PCA) and receiver operating characteristic (ROC) curves to investigate whether data from different groups could be effectively distinguished. Prior to PCA, we standardized the posteriors $\theta$ as $\frac{\theta-\mu}{\sigma}$, where $\mu$ is the mean and $\sigma$ is the standard deviation. Then, we processed 40 parameters with a 16-dimensional posterior parameter set through PCA to reduce them to a single-dimensional component using the scikit-learn library. The reduced component observations corresponding to a specific site were grouped in pairs and represented via ROC curves. A perfect separation between two groups has an area under the curve (AUC) of 1, while an AUC value around 0.5 indicates minimal separation. The results given in Figure  13 suggest that the posterior parameters are not well separated between the distal vs. middle (AUC = 0.50), the middle vs. proximal (AUC = 0.51) and the distal vs. proximal (AUC = 0.51) sites. Furthermore, the kernel density distribution exhibited comparable findings, suggesting a lack of clear separation (Figure 13). Finally, using stochastically inferred parameters from in-silico simulations, it is also possible to investigate the stress uncertainties of the fiber and matrix constituents. A typical stress band observed for the distal site is represented in Figure 14, and for middle and proximal results are given in Supplementary Figure S6.
@@ 4 Discussion and Conclusions
In-silico simulations using computational techniques have significant potential to model very specific cases, ranging from patient-specific tissue behavior to a large variation in population. One significant factor that influences the outcomes of the simulation is the selected constitutive material model, as well as the parameters used in this model. Homogeneous isotropic hyperelastic materials can capture the biomechanical response of many deformable solids. However, most biological materials exhibit inherent variation in their response and may not be accurately represented by a constitutive model with a single set of deterministic parameters. In this approach, the model is typically verified through forward model evaluations using a certain number of experiments. For the backward problem, an inverse approach is mostly used to identify the input parameters from the output. This method is powerful for the characterization of patient- or specimen-specific materials. To account for the uncertainties in the experimental observations, the standard Bayesian approach is widely implemented in many studies [madireddy_bayesian_2015, teferra_bayesian_2019, brewick_uncertainty_2018, staber_stochastic_2017, mihai_LA]. However, it may not always be possible to derive and express the likelihood function in explicit form, particularly for complex loading scenarios. We suggest a likelihood-free approximate Bayesian computation based on a black-box simulator. We demonstrate its implementation for Achilles tendon specimens from diabetic patients, which include deep uncertainties in their in-vitro observations.
@@
There are a wide variety of factors, including pathology, microstructural characteristics, tissue age, and embalming duration, that can significantly influence the biomechanical response. The goal is to stochastically calibrate parameters that represent the uncertain bounds of experimental observations without using the likelihood function. We consider the FRVE model for the representation of in-vitro simulations with 16 parameters. We employ ABC-free likelihood coupled with in-silico simulation and in-vitro measurement for the estimation of posterior parameters. Additionally, Sobol sensitivity analysis and feature importance through a random forest are performed to investigate the influence of parameters on the model. The overall results we observed are as follows. First, in vitro measurements exhibit extreme uncertainties and variability between samples is high. Second, Bayesian optimization converges rapidly with just a few initial iterations and calibrates 16 specimen-specific parameters. Third, sensitivity analysis and feature importance showed that $E_{f}$, $\alpha_{f}$, $g_{1m}$, and $E_{m}$ parameters have the highest influence on the model. Additionally, repeated measures correlation showed a significant correlation between the matrix and fiber viscoelastic parameters. Fourth, the results obtained from ABC demonstrated that the stochastically inferred model parameters can capture the uncertainty band of in-vitro measurements ($y_{obs}$). Fifth, the PCA results demonstrated that the posterior parameters are not separated between sites. For clinical implementation, this observation may suggest that treatment or rehabilitation strategies regarding the mechanical response of diabetic tendons may be generalized to target the entire tendon rather than focusing on specific sites. However, additional data are needed to investigate whether these findings hold in different patient populations.
@@
This study shows that the parameters calibrated using the probabilistic technique (Table 3) deviate from those obtained through the frequentist approach (Table 2). In the frequentist method, the parameter of each specimen is estimated independently based solely on BO. In contrast, the ABC approach incorporates prior knowledge and updates the parameter estimates probabilistically. For example, consider the elastic modulus of fiber $E_{f}$ for the distal site; using the frequentist approach, the average $E_{f}$ value for nine specimens is 204.82 ± 27.42 MPa (mean ± standard deviation). However, when calibrated using ABC, the estimated $E_{f}$ is 216.2 ± 31.12 MPa. This difference highlights a fundamental distinction between the two approaches. However, the following limitations should be addressed in future studies. First, the number of samples is important for the quantification of uncertainty; hence, more specimens should be used to investigate the posterior parameters of the population. Second, investigations are performed on samples with approximately similar geometrical dimensions. Further studies can include geometrical uncertainty. Third, while the concept of ABC is simple and widely applicable, it is computationally expensive as it requires simulating the forward model many times. Additional studies can explore the use of surrogate models based on machine or deep learning algorithms.
In conclusion, the likelihood-free uncertainty quantification approach provides a useful framework for the stochastic calibration of constitutive material model parameters without the need to derive a likelihood function. This general framework can be applied to various combinations of loading data and constitutive models. In medicine, microscale characterization and uncertainty quantification of constituents that cannot be easily determined using experimental techniques, including constituent parameters, constituent stress, and their correlations, could help create more effective therapeutic strategies.
[…]

### v2

@@ 3.1 Bayesian optimization
We first performed specimen-specific parameter calibration using Bayesian optimization (BO), and the results observed for a typical specimen are represented in Figures 4, 5, 6, and 7. Figure 4 shows that the minimum $f(x)$ function converges rapidly, even with only a few initial iterations. This characteristic makes BO popular in machine learning for hyperparameter optimization, where the learning algorithm is treated as a black-box function. Consider the convergence plot given in Figure 4; although the initial minimum value $f(x)$ is 65 N, after 100 iterations (or function calls), it decreases to less than 20 N. In addition, Figure 5 presents the in-vitro measurements obtained from experiments and the in-silico observations generated using optimized FRVE parameter sets in BO. Here, we note that the terms ‘ground truth’, ‘in-vitro measurements’, and ‘experimental observations’ are often used interchangeably to convey a similar meaning. Furthermore, ’in-silico simulation’, ’black-box simulator’, and ’numerical observations’ are also often used interchangeably to refer to a similar meaning. Figure 5 shows that the minimum $f(x)$ values of the objective function, determined by incremental cyclic loading test measurements in one step, two steps, and incremental steps, are 5.693 N, 3.871 N, and 13.232 N, respectively. These results show that the FRVE model can capture the mechanical response of the tendon.
We performed the BO process for 100 calls of the black-box simulator, and the parameter sensitivity can probably be improved once the number of iterations is increased. However, this can increase computational time, and we have not focused on characterizing the parameters with a high rank of sensitivity. The target of using BO is to determine how the specimen-specific parameters vary across specimens and to use these parameters to predict the priors for ABC inference. We present some parameters for the pairs of matrix and fiber constituents that were sampled during BO updating as scatter (Figure 6) and contour (Figure 7) plots to visualize where they are mainly clustered. For example, consider the sample points of $E_{m}$ and $E_{f}$; the elastic modulus of the matrix and fiber components is predominantly updated to around 60 MPa and 240 MPa, respectively. The remaining parameters for a typical specimen-specific calibration are given as additional figures (See Supplementary Figures S1 and S2). We then automated the process to characterize the remaining 26 samples, and the overall results are given in Table 2. These results suggest that inter-subject or inter-specimen variability and uncertainty in the biomechanical properties of these materials are high.
@@
[…]
@@
For the clinical aspect, correlation analysis between micro-scale matrix and fiber parameters calibrated via BO can also be performed. Here, to account for dependency due to repeated observations for the same patient’s specimen harvested from distal, middle, and proximal sites, we used the rm_corr() repeated measures correlation function from the Pingouin statistical package. Repeated measures reveal that the most significant correlations are observed between $\tau_{3m}$ and $\tau_{1f}$ (r = 0.8), $g_{3m}$ and $g_{3f}$ (r = 0.71), $\tau_{2m}$ and $\tau_{1f}$ (r = 0.69), $g_{2m}$ and $\tau_{1f}$ (r = 0.65), $g_{2m}$ and $g_{3f}$ (r = 0.64), while the only negative significant correlation exists between $E_{m}$ and $\tau_{1f}$ (r = 0.48) (p << 0.05) (Figure 8).
Figure 8: Correlation levels between matrix and fiber variables. *Repeated measures correlation coefficient (r) is significant at the p << 0.05 level
@@ 3.2 Sensitivity analysis
To determine the influence of each parameter on the objective function of the FRVE model parameters, we investigate the percentage importance of the parameters using random forest and their sensitivity using Sobol. Here, we use the observations of the error function $\min f(x)$ given in Eq. 9 as output during sensitivity analysis and importance identification. The results of the random forest for a typical specimen in three loading conditions are represented in Figure 9. As expected, the elastic modulus of the fiber $E_{f}$, the exponential coefficient $\alpha_{f}$, and the elastic modulus of the matrix $E_{m}$ have the highest feature importance. Interestingly, the viscoelastic characteristics $g_{im}$ and relaxation time coefficients $\tau_{im}$ of the ground matrix at the $i$th Prony order are more important than the fiber’s parameters for the combined three-loading $\min f(x)$ function.
@@
For Sobol’s sensitivity analysis, the parameter space is created using Saltelli sampling. We generate 20,000 input sampling sets, each consisting of 16 parameters (8 for the ground matrix and 8 for the fiber), and collect the output $\min f(x)$ observations. Similar to the results obtained from the random forest analysis, $E_{f}$, $E_{m}$, and $\alpha_{f}$ are identified as the most sensitive parameters in the model. The total and first-order sensitivity indices for these parameters are presented in Figure 10. Based on the total-order index scores, the three most sensitive parameters, in descending order, are: $E_{f}$ (0.7325), $E_{m}$ (0.4905), and $\alpha_{f}$ (0.3587).
@@ 3.3 Approximate Bayesian computation
We presented specimen-specific parameter identification results in the previous section and observed that the variability in mechanical parameters across samples is high, potentially influenced by factors such as donor age, genetics, tissue pathology, gender, and other sample-specific characteristics. To determine the uncertainty in these parameters and stochastically calibrate them, we implement likelihood-free inference in a probabilistic framework. The ABC inference is initiated with a tolerance threshold of $\epsilon=0.3$, selected empirically to allow for sufficient sampling diversity in the initial generation. This threshold is adaptively updated by the ABC-SMC algorithm in subsequent populations, progressively tightening the acceptance criterion around the observed data. In each generation, 40 accepted sample sets are retained.The inferred parameters for the distal, middle, and proximal sites are given in Table 3. Consider $E_{f}$ at the distal site: the mean is 216.2 MPa, and the 95% HDI bounds are 164.6 MPa and 267.8 MPa.
@@
We visualized the posterior joint distributions to investigate the relationships, correlations, and uncertainty between the parameters after updating beliefs based on in vitro measurement observations. Some parameter pairs plotted for the matrix and fiber components, as well as their distributions for the distal site, are shown in Figure 11. Consider $g_{1}$: there is a significant negative correlation between the matrix and the fiber component (r = -0.37, p << 0.05). However, the remaining parameters shown in Figure 11 for the matrix ($\tau_{1m}$, $g_{2m}$, $E_{m}$, $v_{m}$) and the fiber ($\tau_{1f}$, $g_{1f}$, $g_{2f}$, $\alpha_{f}$) are not significantly correlated. The joint posterior plots inferred for 16 parameters for three sites are given in Supplementary Figures S3-S5. Once the calibration of the parameters is complete and their uncertain ranges for the highest density interval (HDI) 95% are inferred, the next step is to investigate how well the simulated data from the posterior $f(y_{sim}|\theta)$ can capture the uncertainties of in vitro observations $y_{obs}$. We perform forward simulations based on posterior parameters ($\theta$) for two cases. In the first case, we performed three simulations: one based on the mean and the other two based on the low and high bounds of $\theta$ given in Table 3. We then collected observations from $y_{sim}$. The inferred observation band is shown in cyan, and its bounds are represented in blue in Figure 12. In the second case, we performed 40 simulations, each corresponding to a single set of posterior parameters, and then computed the 95% confidence interval (CI) of the 40 $y_{sim}$ observations. The observation band for the 95% CI and its bounds are colored in red (Figure 12). The results obtained in both cases demonstrate that the stochastically calibrated model parameters can capture the uncertainty band of in vitro measurements ($y_{obs}$). However, to characterize and simulate the soft tissue response under extreme uncertainty, it is preferable to use the 95% HDI bound of $\theta$, as demonstrated in case 1.
@@
In some clinical biomechanical studies, it may be necessary to compare the data obtained from different groups. This is usually carried out by applying parametric or non-parametric statistical tests to the results. This approach is effective within a frequentist framework. Although this study does not directly investigate the differences between the three sites, we performed unsupervised machine learning to determine whether the probabilistic data of the groups are clustered and separated. Here, we primarily used principal component analysis (PCA) and receiver operating characteristic (ROC) curves to investigate whether data from different groups could be effectively distinguished. Prior to PCA, we standardized the posteriors $\theta$ as $\frac{\theta-\mu}{\sigma}$, where $\mu$ is the mean and $\sigma$ is the standard deviation. Then, we processed 40 parameters with a 16-dimensional posterior parameter set through PCA to reduce them to a single-dimensional component using the scikit-learn library. The reduced component observations corresponding to a specific site were grouped in pairs and represented via ROC curves. A perfect separation between two groups has an area under the curve (AUC) of 1, while an AUC value around 0.5 indicates minimal separation. The results given in Figure  13 suggest that the posterior parameters are not well separated between the distal vs. middle sites (AUC = 0.50), the middle vs. proximal (AUC = 0.51) and the distal vs. proximal sites (AUC = 0.51). Furthermore, the kernel density distribution exhibited comparable findings, suggesting a lack of clear separation (Figure 13). Finally, using stochastically inferred parameters from in-silico simulations, it is also possible to investigate the stress uncertainties of the fiber and matrix constituents. A typical stress band observed for the distal site is represented in Figure 14, and for the middle and proximal results are given in Supplementary Figure S6.
@@ 4 Discussion and Conclusions
In-silico simulations using computational techniques have significant potential to model very specific cases, ranging from patient-specific tissue behavior to a large variation in population. One significant factor that influences the outcomes of the simulation is the selected constitutive material model, as well as the parameters used in this model. Homogeneous isotropic hyperelastic materials can capture the biomechanical response of many deformable solids. However, most biological materials exhibit an inherent variation in their response and may not be accurately represented by a constitutive model with a single set of deterministic parameters. In this approach, the model is typically verified through forward model evaluations using a certain number of experiments. For the backward problem, an inverse approach is mostly used to identify the input parameters from the output. This method is powerful for the characterization of patient- or specimen-specific materials. To account for the uncertainties in the experimental observations, the standard Bayesian approach is widely implemented in many studies [15, 16, 17, 18, 13]. However, it may not always be possible to derive and express the likelihood function in explicit form, particularly for complex loading scenarios. We suggest a likelihood-free approximate Bayesian computation based on a black-box simulator. We demonstrate its implementation for Achilles tendon specimens from diabetic patients, which include deep uncertainties in their in-vitro observations.
@@
There are a wide variety of factors, including pathology, microstructural characteristics, tissue age, and embalming duration, that can significantly influence the biomechanical response. The goal is to stochastically calibrate parameters that represent the uncertain bounds of experimental observations without using the likelihood function. We consider the FRVE model for the representation of in-vitro simulations with 16 parameters. Sobol sensitivity analysis and feature importance through a random forest are performed to investigate the influence of parameters on the model. The overall results we observed are as follows. First, in vitro measurements exhibit extreme uncertainties and variability between samples is high. Second, Bayesian optimization converges rapidly with just a few initial iterations and calibrates 16 specimen-specific parameters. Third, sensitivity analysis and feature importance showed that $E_{f}$, $\alpha_{f}$, and $E_{m}$ parameters have the highest influence on the model. Additionally, repeated measures correlation showed a significant correlation between the matrix and fiber viscoelastic parameters. Fourth, the results obtained from ABC demonstrated that the stochastically inferred model parameters can capture the uncertainty band of in-vitro measurements ($y_{obs}$). Fifth, the PCA results demonstrated that the posterior parameters are not separated between sites. For clinical implementation, this observation may suggest that treatment or rehabilitation strategies regarding the mechanical response of diabetic tendons may be generalized to target the entire tendon rather than focusing on specific sites. However, additional data are needed to investigate whether these findings hold in different patient populations.
@@
This study shows that the parameters calibrated using the probabilistic technique (Table 3) differ from those obtained through the frequentist approach (Table 2). In the frequentist approach, we optimize parameters separately for each specimen based on its own data using Bayesian Optimization (BO). In contrast, in the ABC approach, we use data from all specimens collectively to infer posterior parameter distributions, capturing uncertainty in the estimates. For example, consider the elastic modulus of fiber $E_{f}$ for the distal site: using the frequentist approach, the average $E_{f}$ value for nine specimens is 204.82 ± 27.42 MPa (mean ± standard deviation), whereas the ABC estimate is 216.2 ± 31.12 MPa. While this difference illustrates how inference outcomes may vary between methods, it should be interpreted with caution due to key modeling assumptions. For instance, the initial choice of $\epsilon=0.3$ was made empirically. While the ABC-SMC algorithm adaptively updates this threshold in later generations, the initial value still reflects a modeling decision that may influence the resulting posterior distributions. Furthermore, the following limitations should be addressed in future studies. First, the number of samples is important for the quantification of uncertainty; hence, more specimens should be used to investigate the posterior parameters of the population. Second, investigations are performed on samples with approximately similar geometrical dimensions. Further studies can include geometrical uncertainty. Third, while the concept of ABC is simple and widely applicable, it can be computationally expensive, as it requires repeatedly simulating the forward model (e.g., using the finite element method). Future studies could explore the use of surrogate models based on machine learning or deep learning algorithms to reduce the computational cost of the forward simulator. Finally, this study focused solely on quantifying uncertainty arising from variability in mechanical parameters across specimens with approximately similar geometry. Other important sources of uncertainty—such as measurement noise, model-form errors, and geometric variability—were not considered in the current framework. Future work could extend this approach by incorporating these additional sources of uncertainty.
In conclusion, the likelihood-free uncertainty quantification approach provides a useful framework for the stochastic calibration of constitutive material model parameters without the need to derive a likelihood function. This general framework can be applied to various combinations of loading data and constitutive models. In medicine, microscale characterization and quantification of uncertainty of constituents that cannot be easily determined using experimental techniques, including constituent parameters, constituent stress, and their correlations, could help create more effective therapeutic strategies.
5 Acknowledgments
The author thanks Prof. Firat Ozan from Kayseri City Hospital for his constructive comments and suggestions for this work.